# From Text to Multimodality: Exploring the Evolution and Impact of Large Language Models in Medical Practice

### v5

@@ Abstract
Index Terms:
large language models, medical practice, multimodality, artificial intelligence
Index Terms:
Multimodal Large Language Models (MLLMs), Medical Imaging, Clinical Decision Support, Patient Engagement, Data Integration
@@ I Introduction
The landscape of healthcare is constantly evolving, driven by an unprecedented explosion of data. Electronic health records, medical imaging, genomic sequencing, and wearable sensors generate an overwhelming amount of information, exceeding human capacity for efficient analysis and interpretation [1]. This phenomenon presents both an opportunity and a challenge: ingesting this information can revolutionize healthcare, but doing so requires innovative tools capable of processing and synthesizing these diverse data streams. Artificial intelligence (AI) has emerged as a powerful force in addressing this challenge, with large language models (LLMs) at the forefront of this revolution.
@@
Initially, LLMs focused primarily on text-based tasks, demonstrating remarkable proficiency in understanding and generating human-like language [2]. However, the inherent multimodality of medicine, where clinical decisions often rely on the synthesis of information from diverse sources such as images, text, and genomics, necessitates more versatile models [3]. This need has given rise to Multimodal Large Language Models (MLLMs), a new generation of LLMs capable of processing and integrating information from various modalities. These advanced models potentially unlock a new era of precision medicine and personalized healthcare, offering a more comprehensive approach to medical data analysis and decision-making.
@@
A key strength of MLLMs is their ability to bridge the gap between unstructured and structured data, a particularly valuable feature in healthcare where information is often fragmented across different formats. For example, the REALM framework leverages LLMs to encode clinical notes and integrates them with time-series EHR data, enhancing clinical predictions by incorporating external knowledge from knowledge graphs [4]. In a similar vein, the MedDr model [5] employs a diagnosis-guided bootstrapping strategy to build vision-language datasets, showcasing superior performance across various medical tasks through a retrieval-augmented diagnosis approach. These advancements underscore the potential of MLLMs to enhance data interoperability and extract meaningful insights from diverse sources, potentially revolutionizing how healthcare professionals access and utilize patient information.
@@
MLLMs show great potential for transforming healthcare by enabling a more comprehensive understanding of patient health, potentially leading to improved diagnostics, personalized treatment plans, and enhanced patient engagement [6]. For instance, these models could analyze a patient’s medical history, imaging scans, and genetic data to provide more accurate diagnoses and predict disease risks, facilitating early intervention and tailored treatment strategies. In the field of medical imaging, the integration of LLMs has demonstrated significant progress. Research has shown the effectiveness of visual language models (VLMs), a subset of MLLMs, in analyzing various biomedical images, including brain MRIs, blood cell images, and chest X-rays [7]. A notable example is the LlaVA-Rad model, a lightweight and open-source multimodal system that has achieved state-of-the-art results on standard radiology tasks. This model has surpassed larger counterparts in both performance and accessibility, making it particularly suitable for real-world clinical applications [6].
@@
MLLMs could also enhance communication between patients and healthcare providers through interactive chatbots and virtual assistants, potentially improving patient engagement and healthcare accessibility [8]. The creation of chatbots like MedAide, which utilize optimized tiny-LLMs on edge devices, demonstrates the capacity of MLLMs to provide medical assistance in resource-limited settings and remote areas, addressing challenges in healthcare access [9]. However, developing reliable and trustworthy medical chatbots requires addressing critical issues such as accuracy, privacy protection, and bias mitigation to meet the high standards required for patient care and safety.
@@
Our review aims to offer an overview of the current state of MLLMs in medicine and healthcare. We will not only examine their architecture, capabilities, and limitations, but also explore potential applications across various medical domains. We will critically assess the challenges and research gaps impeding the widespread adoption of MLLMs in clinical settings, including data limitations, technical difficulties, and ethical considerations [10]. For example, the evaluation of LLMs in healthcare often relies on benchmarks that are usually unfit for real-world diagnostic frameworks and are likely vulnerable to data leakage [11], which indicates the need for standardized evaluation frameworks and comprehensive datasets that accurately reflect the clinical practice. By analyzing the current research landscape and identifying key areas for further development, this review seeks to guide the responsible and effective integration of MLLMs into healthcare. Our goal is to contribute to a brighter future where AI assists clinicians and enhances patient care, while addressing the unique challenges and requirements within the field. In order to provide a clear overview of the various applications and components of MLLMs in medicine, we present a taxonomy Fig 1. This simplified taxonomy categorizes the key aspects of MLLMs in healthcare and medicine.
{forest}
@@
in Medicine, root, [Applications in
Medicine, level1, [Clinical Decision Support, level2, [Diagnosis and treatment
recommendations [1, 12], leaf3] [Prognostic prediction and
risk stratification [13], leaf3] ] [Medical Imaging, level2, [Integration of visual and
textual data [14], leaf4] [Report generation and
VQA [15, 16], leaf4] ] [Patient Engagement and
Communication, level2, [Chatbots & virtual assistants [9], leaf5] [Personalized health
recommendations [17], leaf5] [Patient education [18], leaf5] ] ] [Technical Aspects, level1, [Modality Alignment
Methods, level2, [Multimodal Converters:
Images to text [19], leaf6] [Multimodal Perceivers:
Vision transformers [15], leaf6] [Tools Assistance: Knowledge
graphs [20, 21], leaf6] [Data-Driven Methods: Using
large datasets [22, 23], leaf6] ] [Knowledge Integration, level2, [Format unification [24, 25], leaf7] [Graph-based methods [26, 27], leaf7] ] ] [Challenges and
Limitations, level1, [Data Scarcity and Quality, level2, [Limited large-scale
multimodal datasets [22], leaf1] [Biases in training data [28], leaf1] ] [Ethical and Regulatory
Considerations, level2, [Privacy & data security [29], leaf2] [Bias and fairness in
clinical outcomes [10, 13, 30], leaf2] ] [Model Interpretability and
Explainability, level2, [Trust and transparency in
clinical decisions [3], leaf3] [Overcoming the
”black box” nature [31], leaf3] ] ] [Future Directions, level1, [Data Augmentation and
Multilingual Support, level2, [Development of large
multilingual datasets [32], leaf4] [Privacy-preserving
data sharing [33], leaf4] ] [Advancements in
Modality Alignment, level2, [Methods for cross-modality
data representation [34], leaf5] ] [Ethics and Compliance, level2, [Regulatory frameworks specific
to MLLMs [35], leaf6] ] ] [Key Research
Gaps, level1, [Dataset standardization, leaf7] [Ethical guidelines for
multimodal AI systems, leaf7] [Improvements in modality
alignment and generalization, leaf7] ] ]
@@ II-A LLMs and Their Evolution: From Text to Multimodal Understanding
Large language models (LLMs) represent a significant advancement in artificial intelligence, demonstrating remarkable capabilities in comprehending and generating human-like text. Architecturally, they often rely on the Transformer network [36], a powerful neural network structure that excels at capturing long-range dependencies and contextual relationships within text. LLMs are initially trained on massive text corpora, a process known as pre-training, to develop a generalized understanding of language structure and patterns. This pre-training phase allows them to learn a wide range of linguistic features and relationships, making them adaptable to various downstream tasks.
LLMs can be fine-tuned on smaller, task-specific datasets to further refine their performance in specific domains. For example, ClinicalT5 [37] demonstrates how a general-purpose LLM (T5) can be adapted for clinical text by fine-tuning it on the MIMIC-III dataset. This adaptation to the medical domain is crucial for addressing the unique challenges of medical language, including its specialized vocabulary and complex semantic relationships [38].
Despite impressive capabilities, LLMs may face many limitations. One notable issue is ”hallucination”, where the model generates plausible but incorrect or nonsensical information, as highlighted in the study by Ziaei and Schmidgall [39]. Hallucination can be particularly problematic in healthcare, where accuracy and reliability are the top priorities [28, 40]. Additionally, biases present in the training data can propagate to the model’s outputs, leading to unfair or discriminatory outcomes, as discussed in the paper by Reddy [41]. Addressing these biases requires careful data curation and model development strategies [42].
Integrating LLMs with other modalities, such as images and videos, results in MLLMs. MLLMs like GPT-4V [43, 44] and Gemini [45, 46] process and generate both text and visual information, which opens up new possibilities for medical applications. For example, MLLMs can be used to generate captions for medical images [47], answer visual questions about medical images [48], and even assist in medical report generation [49]. On the other hand, MLLMs are still in their early stages, and these models often face challenges in terms of accuracy, reliability, and ethical considerations [50]. Further research is needed to fully explore the potential of MLLMs and address these challenges to enable their safe and effective deployment in clinical practice.
@@ II-B Multimodality in Medicine: Embracing the Rich Variety of Data
Medicine is inherently multimodal, as it involves many types of information beyond just written text. For example, when a patient comes in with a possible lung infection, their case might include several kinds of data: written information like their medical history and symptoms noted by doctors, images from chest X-rays, sound recordings of their breathing, and even genetic information to assess their personal risk. Combining these different types of information is important for getting a complete picture of a patient’s health and more accurate and personalized medical care [51]. This is where multimodal models shine, because they are designed to process and integrate various types of data.
We have seen a surge in developing MLLMs capable of processing and integrating diverse medical data types [52]. The study by Tian et al. [53] exemplifies this trend, showcasing a Med-MLLM model that handles both visual and textual data for improved clinical decision-making, particularly in rare diseases and pandemics. MLLMs could revolutionize various medical practices. For instance, in radiology, MLLMs are being explored for generating comprehensive reports [54], assisting in diagnosis by analyzing both images and clinical notes [4], and facilitating visual search and querying within patient imaging history [55].
MLLMs for more specialized medical tasks has also gained momentum. SkinGPT-4 is a system designed for dermatological diagnosis using both images and clinical data, which offers autonomous image evaluation and treatment recommendations [16]. Developing robust and reliable MLLMs requires overcoming many challenges. Large, diverse, and unbiased medical datasets across multiple modalities are crucial [48]. Accuracy, interpretability, explainability, interoperability, and ethics are important to be discussed before integrating into existing clinical workflows [35].
@@ II-C Modality Alignment Methods: Bridging the Semantic Gap
  • Multimodal Converters: These methods transform data from different modalities into a unified representation that LLMs can understand. For example, images might be converted into textual descriptions or embeddings before being fed into the LLM. This approach is seen in models like X-LLM [19], which treats modalities as foreign languages and converts them to text, or LIFTED [56], which transforms modalities into natural language descriptions for improved clinical trial outcome prediction.
  • Multimodal Perceivers: These methods directly enhance the LLMs’ perception of multimodal data. A vision encoder can be integrated into the LLM architecture to enable it to directly process and understand images and texts. Med-Flamingo [15] incorporates a vision transformer for medical image understanding. Similar approaches can be seen in models like SkinGPT-4 [16], which combines a vision transformer with a LLM for dermatological diagnosis, and MedVersa [34], which uses a LLM as a learnable orchestrator to process both visual and linguistic information to interpret medical images.
  • Tools Assistance: These methods uses external tools for multimodalities. A knowledge graph can link concepts across modalities and provide additional context for the LLM. The study by Gao et al. [20] uses the Unified Medical Language System (UMLS) knowledge to enhance diagnosis generation. A similar approach is used in BioLORD-2023 [21], which integrates LLMs with knowledge graphs to improve performance in semantic textual similarity, biomedical concept representation, and named entity linking.
  • Data-Driven Methods: These methods rely on large-scale multimodal datasets to train LLMs directly on multimodal tasks and often involves developing new architectures and training strategies so the modals can learn from different modalities simultaneously. Models like BiomedGPT [23] are trained on those diverse multimodal datasets. The recent open-source frameworks like Hippocrates [22] further facilitates this approach by providing access to training datasets, codebases, and evaluation protocols, encouraging further collaborative efforts.
Each method has its own strengths and weaknesses. Multimodal converters are relatively simple but may lead to information loss during conversion [57]. Multimodal perceivers can potentially capture richer relationships between modalities, but requires more complex architectures and training. Tools assistance uses existing knowledge bases and resources but may not be comprehensive or up-to-date. Data-driven methods can achieve high performance but require large and diverse datasets.
@@ III Applications of MLLMs in Medicine
While MLLMs integrate diverse data modalities, offering a more comprehensive view of patient health and the ability to detect complex patterns for improved diagnosis, treatment personalization, and risk assessment [58], their development is still in the early stages. As a result, LLMs continue to dominate the field due to their maturity and established performance. This section introduces both LLMs and MLLMs, while emphasizing the promise of multimodal models.
@@
Diagnosis and Treatment Recommendations: NYUTron is an LLM trained on clinical notes, for predicting patient outcomes with high accuracy [1]. PMC-LLaMA is a performant LLM for medical Q&A [59, 60, 61]. Almanac is an LLM augmented with retrieval capabilities from curated medical resources and has significant improvements in factuality, completeness, user preference, and safety for clinical decision-making [12]. Med-PaLM 2 is a specialized LLM for medicine, showcased superior performance on medical question answering and treatment recommendation tasks, significantly outperforming GPT-3.5 [2]. Med-PaLM M is a MLLM achieving competitive performance on medical question answering, radiology report generation, etc. [62].
@@
Prognostic Prediction and Risk Stratification: Beyond diagnosis, LLMs have also shown promise in prognostic prediction and risk stratification. Researchers have tried to use LLMs for prognostic prediction in immunotherapy, achieving encouraging results in improving accuracy and facilitating early disease detection [13]. Studies have demonstrated the potential of LLMs and MLLMs to predict outcomes like mortality, length of stay, and readmission using structured EHR data, outperforming traditional machine learning models in few-shot settings. [63, 26] The Health-LLM, a LLM framework that has vision capability in the future integrating health reports and medical knowledge into LLMs, has also been proposed for enhanced disease prediction and personalized health management, showcasing its superior performance over existing systems [64].
@@
Despite the potentials, LLM and MLLMs face limitations in clinical decision support. Explainability and interpretability remain challenging, as their complex decision-making processes often lack transparency, hindering clinician trust and understanding [65]. Another concern is the potential for bias and unfairness due to inherent biases in MLLM training data, which can exacerbate healthcare disparities [66]. Extensive real-world validation in diverse clinical settings is crucial to ensure the effectiveness and safety of MLLMs before widespread adoption, addressing potential risks and unexpected outcomes [35].
@@
Several real-world case studies have demonstrated the potential of LLMs in clinical decision support. One study showed that an LLM optimized for diagnostic reasoning improved clinicians’ differential diagnosis accuracy on challenging medical cases [67]. Another study found that an LLM could accurately classify patient acuity levels in the emergency department, comparable to human physicians [31].
@@
The performance of LLMs in clinical decision support is often evaluated using traditional metrics like accuracy, precision, recall, F1 score, and AUC [68]. Evaluating their effectiveness in this context requires moving beyond accuracy and considering additional factors like interpretability, fairness, impact on clinical workflows, and user trust [69]. The development of standardized evaluation frameworks and benchmarks, such as CLUE and BenchHealth, is crucial for assessing the clinical performance and real-world applicability of LLMs [70, 71, 72].

### v6

@@ I Introduction
The landscape of healthcare is constantly evolving, driven by an unprecedented explosion of data. Electronic health records, medical imaging, genomic sequencing, and wearable sensors generate an overwhelming amount of information, exceeding human capacity for efficient analysis and interpretation [Jiang2023Health]. This phenomenon presents both an opportunity and a challenge: ingesting this information can revolutionize healthcare, but doing so requires innovative tools capable of processing and synthesizing these diverse data streams. Artificial intelligence (AI) has emerged as a powerful force in addressing this challenge, with large language models (LLMs) at the forefront of this revolution.
@@
Initially, LLMs focused primarily on text-based tasks, demonstrating remarkable proficiency in understanding and generating human-like language [Singhal2023Large]. However, the inherent multimodality of medicine, where clinical decisions often rely on the synthesis of information from diverse sources such as images, text, and genomics, necessitates more versatile models [Acosta2022Multimodal]. This need has given rise to Multimodal Large Language Models (MLLMs), a new generation of LLMs capable of processing and integrating information from various modalities. These advanced models potentially unlock a new era of precision medicine and personalized healthcare, offering a more comprehensive approach to medical data analysis and decision-making.
@@
A key strength of MLLMs is their ability to bridge the gap between unstructured and structured data, a particularly valuable feature in healthcare where information is often fragmented across different formats. For example, the REALM framework leverages LLMs to encode clinical notes and integrates them with time-series EHR data, enhancing clinical predictions by incorporating external knowledge from knowledge graphs [Yinghao2024REALM]. In a similar vein, the MedDr model [He2024MedDr] employs a diagnosis-guided bootstrapping strategy to build vision-language datasets, showcasing superior performance across various medical tasks through a retrieval-augmented diagnosis approach. These advancements underscore the potential of MLLMs to enhance data interoperability and extract meaningful insights from diverse sources, potentially revolutionizing how healthcare professionals access and utilize patient information.
@@
MLLMs show great potential for transforming healthcare by enabling a more comprehensive understanding of patient health, potentially leading to improved diagnostics, personalized treatment plans, and enhanced patient engagement [chaves2024training]. For instance, these models could analyze a patient’s medical history, imaging scans, and genetic data to provide more accurate diagnoses and predict disease risks, facilitating early intervention and tailored treatment strategies. In the field of medical imaging, the integration of LLMs has demonstrated significant progress. Research has shown the effectiveness of visual language models (VLMs), a subset of MLLMs, in analyzing various biomedical images, including brain MRIs, blood cell images, and chest X-rays [Van2024On]. A notable example is the LlaVA-Rad model, a lightweight and open-source multimodal system that has achieved state-of-the-art results on standard radiology tasks. This model has surpassed larger counterparts in both performance and accessibility, making it particularly suitable for real-world clinical applications [chaves2024training].
@@
MLLMs could also enhance communication between patients and healthcare providers through interactive chatbots and virtual assistants, potentially improving patient engagement and healthcare accessibility [Mesko2023Impact]. The creation of chatbots like MedAide, which utilize optimized tiny-LLMs on edge devices, demonstrates the capacity of MLLMs to provide medical assistance in resource-limited settings and remote areas, addressing challenges in healthcare access [Abdul2024MedAide]. However, developing reliable and trustworthy medical chatbots requires addressing critical issues such as accuracy, privacy protection, and bias mitigation to meet the high standards required for patient care and safety.
@@
Our review aims to offer an overview of the current state of MLLMs in medicine and healthcare. We will not only examine their architecture, capabilities, and limitations, but also explore potential applications across various medical domains. We will critically assess the challenges and research gaps impeding the widespread adoption of MLLMs in clinical settings, including data limitations, technical difficulties, and ethical considerations [Harrer2023Attention]. For example, the evaluation of LLMs in healthcare often relies on benchmarks that are usually unfit for real-world diagnostic frameworks and are likely vulnerable to data leakage [Lyu2024Language], which indicates the need for standardized evaluation frameworks and comprehensive datasets that accurately reflect the clinical practice. By analyzing the current research landscape and identifying key areas for further development, this review seeks to guide the responsible and effective integration of MLLMs into healthcare. Our goal is to contribute to a brighter future where AI assists clinicians and enhances patient care, while addressing the unique challenges and requirements within the field. In order to provide a clear overview of the various applications and components of MLLMs in medicine, we present a taxonomy Fig 1. This simplified taxonomy categorizes the key aspects of MLLMs in healthcare and medicine.
@@
[…]
LLMs and MLLMs in Medicine Applications in Medicine Clinical Decision Support Diagnosis and treatment recommendations [Jiang2023Health, Hiesinger2023Almanac] Prognostic prediction and risk stratification [Zekai2023Language] Medical Imaging Integration of visual and textual data [Jiawei2024MISS] Report generation and VQA [Michael2023Med, Zhou2024Pre] Patient Engagement and Communication Chatbots & virtual assistants [Abdul2024MedAide] Personalized health recommendations [dao2024llm] Patient education [Akash2024MedSumm] Technical Aspects Modality Alignment Methods Multimodal Converters: Images to text [Feilong2023X] Multimodal Perceivers: Vision transformers [Michael2023Med] Tools Assistance: Knowledge graphs [gao2023leveraging, Remy2024BioLORD] Data-Driven Methods: Using large datasets [Emre2024Hippocrates, Kai2023BiomedGPT] Knowledge Integration Format unification [shao2024dual, Qinghao2023mPLUG] Graph-based methods [niu2024ehr, ektefaie2023multimodal] Challenges and Limitations Data Scarcity and Quality Limited large-scale multimodal datasets [Emre2024Hippocrates] Biases in training data [Nassiri2024Recent] Ethical and Regulatory Considerations Privacy & data security [peng2024securinglargelanguagemodels] Bias and fairness in clinical outcomes [Harrer2023Attention, Zekai2023Language, Jiang2024754] Model Interpretability and Explainability Trust and transparency in clinical decisions [Acosta2022Multimodal] Overcoming the ”black box” nature [Williams2024Use] Future Directions Data Augmentation and Multilingual Support Development of large multilingual datasets [Iker2024MedMT5] Privacy-preserving data sharing [Wang2023Accelerating] Advancements in Modality Alignment Methods for cross-modality data representation [Zhou2024Generalist] Ethics and Compliance Regulatory frameworks specific to MLLMs [Ong2024Ethical] Key Research Gaps Dataset standardization Ethical guidelines for multimodal AI systems Improvements in modality alignment and generalization
@@ II-A LLMs and Their Evolution: From Text to Multimodal Understanding
Large language models (LLMs) represent a significant advancement in artificial intelligence, demonstrating remarkable capabilities in comprehending and generating human-like text. Architecturally, they often rely on the Transformer network [vaswani2017attention], a powerful neural network structure that excels at capturing long-range dependencies and contextual relationships within text. LLMs are initially trained on massive text corpora, a process known as pre-training, to develop a generalized understanding of language structure and patterns. This pre-training phase allows them to learn a wide range of linguistic features and relationships, making them adaptable to various downstream tasks.
LLMs can be fine-tuned on smaller, task-specific datasets to further refine their performance in specific domains. For example, ClinicalT5 [Lu2022ClinicalT5] demonstrates how a general-purpose LLM (T5) can be adapted for clinical text by fine-tuning it on the MIMIC-III dataset. This adaptation to the medical domain is crucial for addressing the unique challenges of medical language, including its specialized vocabulary and complex semantic relationships [Meng2024application].
Despite impressive capabilities, LLMs may face many limitations. One notable issue is ”hallucination”, where the model generates plausible but incorrect or nonsensical information, as highlighted in the study by Ziaei and Schmidgall [Rojin2023Language]. Hallucination can be particularly problematic in healthcare, where accuracy and reliability are the top priorities [Nassiri2024Recent, peng2024emerging]. Additionally, biases present in the training data can propagate to the model’s outputs, leading to unfair or discriminatory outcomes, as discussed in the paper by Reddy [Reddy2023Evaluating]. Addressing these biases requires careful data curation and model development strategies [han2024medsafetybenchevaluatingimprovingmedical].
Integrating LLMs with other modalities, such as images and videos, results in MLLMs. MLLMs like GPT-4V [Deng2024When, achiam2023gpt] and Gemini [Ankit2024Gemini, team2023gemini] process and generate both text and visual information, which opens up new possibilities for medical applications. For example, MLLMs can be used to generate captions for medical images [Iryna2024Vision], answer visual questions about medical images [Yutao2024OmniMedVQA], and even assist in medical report generation [Nakaura2024impact]. On the other hand, MLLMs are still in their early stages, and these models often face challenges in terms of accuracy, reliability, and ethical considerations [Yihe2024Unbridled]. Further research is needed to fully explore the potential of MLLMs and address these challenges to enable their safe and effective deployment in clinical practice.
@@ II-B Multimodality in Medicine: Embracing the Rich Variety of Data
Medicine is inherently multimodal, as it involves many types of information beyond just written text. For example, when a patient comes in with a possible lung infection, their case might include several kinds of data: written information like their medical history and symptoms noted by doctors, images from chest X-rays, sound recordings of their breathing, and even genetic information to assess their personal risk. Combining these different types of information is important for getting a complete picture of a patient’s health and more accurate and personalized medical care [Tripathi2024Efficient]. This is where multimodal models shine, because they are designed to process and integrate various types of data.
We have seen a surge in developing MLLMs capable of processing and integrating diverse medical data types [Duzhen2024MM]. The study by Tian et al. [Liu2023medical] exemplifies this trend, showcasing a Med-MLLM model that handles both visual and textual data for improved clinical decision-making, particularly in rare diseases and pandemics. MLLMs could revolutionize various medical practices. For instance, in radiology, MLLMs are being explored for generating comprehensive reports [Bhayana2024Chatbots], assisting in diagnosis by analyzing both images and clinical notes [Yinghao2024REALM], and facilitating visual search and querying within patient imaging history [Yildirim2024Multimodal].
MLLMs for more specialized medical tasks has also gained momentum. SkinGPT-4 is a system designed for dermatological diagnosis using both images and clinical data, which offers autonomous image evaluation and treatment recommendations [Zhou2024Pre]. Developing robust and reliable MLLMs requires overcoming many challenges. Large, diverse, and unbiased medical datasets across multiple modalities are crucial [Yutao2024OmniMedVQA]. Accuracy, interpretability, explainability, interoperability, and ethics are important to be discussed before integrating into existing clinical workflows [Ong2024Ethical].
@@ II-C Modality Alignment Methods: Bridging the Semantic Gap
  • Multimodal Converters: These methods transform data from different modalities into a unified representation that LLMs can understand. For example, images might be converted into textual descriptions or embeddings before being fed into the LLM. This approach is seen in models like X-LLM [Feilong2023X], which treats modalities as foreign languages and converts them to text, or LIFTED [Wenhao2024Multimodal], which transforms modalities into natural language descriptions for improved clinical trial outcome prediction.
  • Multimodal Perceivers: These methods directly enhance the LLMs’ perception of multimodal data. A vision encoder can be integrated into the LLM architecture to enable it to directly process and understand images and texts. Med-Flamingo [Michael2023Med] incorporates a vision transformer for medical image understanding. Similar approaches can be seen in models like SkinGPT-4 [Zhou2024Pre], which combines a vision transformer with a LLM for dermatological diagnosis, and MedVersa [Zhou2024Generalist], which uses a LLM as a learnable orchestrator to process both visual and linguistic information to interpret medical images.
  • Tools Assistance: These methods uses external tools for multimodalities. A knowledge graph can link concepts across modalities and provide additional context for the LLM. The study by Gao et al. [gao2023leveraging] uses the Unified Medical Language System (UMLS) knowledge to enhance diagnosis generation. A similar approach is used in BioLORD-2023 [Remy2024BioLORD], which integrates LLMs with knowledge graphs to improve performance in semantic textual similarity, biomedical concept representation, and named entity linking.
  • Data-Driven Methods: These methods rely on large-scale multimodal datasets to train LLMs directly on multimodal tasks and often involves developing new architectures and training strategies so the modals can learn from different modalities simultaneously. Models like BiomedGPT [Kai2023BiomedGPT] are trained on those diverse multimodal datasets. The recent open-source frameworks like Hippocrates [Emre2024Hippocrates] further facilitates this approach by providing access to training datasets, codebases, and evaluation protocols, encouraging further collaborative efforts.
Each method has its own strengths and weaknesses. Multimodal converters are relatively simple but may lead to information loss during conversion [park2023cross]. Multimodal perceivers can potentially capture richer relationships between modalities, but requires more complex architectures and training. Tools assistance uses existing knowledge bases and resources but may not be comprehensive or up-to-date. Data-driven methods can achieve high performance but require large and diverse datasets.
@@ III Applications of MLLMs in Medicine
While MLLMs integrate diverse data modalities, offering a more comprehensive view of patient health and the ability to detect complex patterns for improved diagnosis, treatment personalization, and risk assessment [Tripathi2024Large], their development is still in the early stages. As a result, LLMs continue to dominate the field due to their maturity and established performance. This section introduces both LLMs and MLLMs, while emphasizing the promise of multimodal models.
@@
Diagnosis and Treatment Recommendations: NYUTron is an LLM trained on clinical notes, for predicting patient outcomes with high accuracy [Jiang2023Health]. PMC-LLaMA is a performant LLM for medical Q&A [Chaoyi2023PMC, Chaoyi2023PMC1, Wu2024PMC]. Almanac is an LLM augmented with retrieval capabilities from curated medical resources and has significant improvements in factuality, completeness, user preference, and safety for clinical decision-making [Hiesinger2023Almanac]. Med-PaLM 2 is a specialized LLM for medicine, showcased superior performance on medical question answering and treatment recommendation tasks, significantly outperforming GPT-3.5 [Singhal2023Large]. Med-PaLM M is a MLLM achieving competitive performance on medical question answering, radiology report generation, etc. [Tu2023Towards].
@@
Prognostic Prediction and Risk Stratification: Beyond diagnosis, LLMs have also shown promise in prognostic prediction and risk stratification. Researchers have tried to use LLMs for prognostic prediction in immunotherapy, achieving encouraging results in improving accuracy and facilitating early disease detection [Zekai2023Language]. Studies have demonstrated the potential of LLMs and MLLMs to predict outcomes like mortality, length of stay, and readmission using structured EHR data, outperforming traditional machine learning models in few-shot settings. [Zhu2024Prompting, niu2024ehr] The Health-LLM, a LLM framework that has vision capability in the future integrating health reports and medical knowledge into LLMs, has also been proposed for enhanced disease prediction and personalized health management, showcasing its superior performance over existing systems [Mingyu2024Health].
@@
Despite the potentials, LLM and MLLMs face limitations in clinical decision support. Explainability and interpretability remain challenging, as their complex decision-making processes often lack transparency, hindering clinician trust and understanding [Park2024Assessing]. Another concern is the potential for bias and unfairness due to inherent biases in MLLM training data, which can exacerbate healthcare disparities [Schmidgall2024AgentClinic]. Extensive real-world validation in diverse clinical settings is crucial to ensure the effectiveness and safety of MLLMs before widespread adoption, addressing potential risks and unexpected outcomes [Ong2024Ethical].
@@
Several real-world case studies have demonstrated the potential of LLMs in clinical decision support. One study showed that an LLM optimized for diagnostic reasoning improved clinicians’ differential diagnosis accuracy on challenging medical cases [Daniel2023Towards]. Another study found that an LLM could accurately classify patient acuity levels in the emergency department, comparable to human physicians [Williams2024Use].
@@
The performance of LLMs in clinical decision support is often evaluated using traditional metrics like accuracy, precision, recall, F1 score, and AUC [Junda2024JMLR]. Evaluating their effectiveness in this context requires moving beyond accuracy and considering additional factors like interpretability, fairness, impact on clinical workflows, and user trust [Kilian2024Review]. The development of standardized evaluation frameworks and benchmarks, such as CLUE and BenchHealth, is crucial for assessing the clinical performance and real-world applicability of LLMs [Amin2024CLUE, Liu2024Large, Congyun2024RJUA].